2(x-3)(x+2)^(2)+2(x+2)(x-3)^(3)
Please factorize the above equation.PLEASE SHOW THE STEP.IM CONFUSED.
THANKS

Answer by stanbon(75887)

2(x-3)(x+2)^(2)+2(x+2)(x-3)^(3)
Please factor
-----------
Common Factor: 2(x+2)(x-3)
----
Divide both terms to get the other factor:
---
= 2(x+2)(x-3)[(x+2) + (x-3)^2]
-------------------------------
= 2(x+2)(x-3) [x+2 + x^2-6x+9]
---
= 2(x+2)(x-3)[x^2-5x+11}
=============================
Cheers,
Stan H.
==================